Circular No. 6866

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 980329 D. M. Cole, A. R. Cooray, J. M. Quashnock, D. E. Vanden Berk,

D. Q. Lamb, D. E. Reichart, G. T. Richards, K. Gloria, D. Long, and T. Hoyes, University of Chicago, on behalf of the Astrophysical Research Consortium (ARC), report: “We made near-infrared observations (IAUC 6860) of the BeppoSAX NFI error circle for GRB 980329 (IAUC 6853, 6854) between Apr. 1.167 and 1.277 UT, using the ARC 3.5-m telescope (+ GRIM II) at Apache Point Observatory. We find no detectable object, down to a limiting magnitude of J = 20.9 ± 0.5 (1σ) at the position of the variable radio source VLA J0702+3850 proposed by Taylor

et al. (http://gcn.gsfc.nasa.gov/gcn/gcn3/040.gcn3 and IAUC 6864)

to be the radio afterglow from GRB 980329. The measurement by Man-

nuci et al. (http://gcn.gsfc.nasa.gov/gcn/gcn3/046.gcn3) of an ob-

ject with J = 19.2_{−0.5}^{+0.8} at the position of the radio source on Mar. 29.84– 29.86 implies that this object dimmed by at least ΔJ = 1.7 ± 0.9 over ∼ 2.4 days. J-band images are available in GIF and PostScript formats at

ftp://astro.chicago.edu/pub/astro/GRB/.”

η CARINAE

R. Viotti and P. Grandi, Istituto di Astrofisica Spaziale, CNR, Rome; L. A. Antonelli and S. Rebecchi, BeppoSAX Science Data Center (SDC), Rome; J. M. Muller, BeppoSAX SDC and Space Research Organization of the Netherlands, Utrecht; and M. Villada, Observatorio Astronomico, Cordoba, report: “η Car was observed as a target-of-opportunity with LECS and MECS onboard BeppoSAX over Mar. 18.1–19.2 UT, during egress from the ongoing spectroscopic event (IAUC 6849). A preliminary comparison with the Dec. 1996 BeppoSAX observations revealed a large flux defect between 1.5 and 6 keV, indicating that the event was still in progress at high energies. The Fe-k line flux appears slightly weaker. The variation can be interpreted as due to a larger amount of absorbing matter (∼ 3 times) from a hot 4.7-keV component, with only minor changes of its unabsorbed flux. Simultaneous optical spectroscopy with the 2.15-m telescope of the Complejo Astronomico El Leoncito provided the following emissionline equivalent widths (in nm): [NII] 575.5-nm, 1.0; He I 587.6-nm, 0.9; He I 667.8-nm, 1.1; He I 706.5-nm, 1.6. This is about half of their intensity at maximum, confirming the ending of the deeper part of the event.”

1998 April 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green